A. before addition of any HBr B. after addition of 12.5 mL of HBr C. after addition of 23.0 mL of HBr D. after addition of 25Calculate the pH for each of the cases in the titration of 25.0 mL of 0.100 M pyridine, C5H5N(aq) with 0.100 M HBr(aq) .
Answer #1

CHENKG = 1.78109 pkb = -log Kb = -log (1-7 X10 9) = 8.77 ① stav +H2O - SHa Nit + oH I 01 - CL to * - X * E 0.1-X Kb = Easts WCH_N= Molarity = 0.1M | Hor: Mohrity = 0-1M 1 ② HBr molarity=0-IM volume=25m/=0.0252 volume = 12.5m1=0.012541 volume=23m1 = 0( HBr. Molarity = 0.19 volume=25mi= 0.0252 no. of moles= 0.18 0.025= 0.0025 GHANT HBr - It, NH BY weak base strongacid salt I(E) HBr Molarity = OIM volume = 31m1= 0.0312 no-of moles a 0-031X0-1=0.0031 G-Hant HBr H, NHBY I 0.0025 0.003) el-0.0025-0-00(E) HBr Molarity = OIM volume = 31m1= 0.0312 no-of moles a 0-031X0-1=0.0031 G-Hant HBr H, NHBY I 0.0025 0.003) el-0.0025-0-00pH formulas for salts weak acid + strong barse -salft ito pH = ± [pkw tpka + loge] strongacid & weakbase salf tito pH = ₃ [